# Eigenstate solutions of the Fermi-Hubbard model via symmetry-enhanced variational quantum eigensolver

### v1

@@ Abstract
The Variational Quantum Eigensolver (VQE), as a hybrid quantum-classical algorithm, is an important tool for effective quantum computing in the current noisy intermediate-scale quantum (NISQ) era. However, the traditional hardware-efficient ansatz without taking into account symmetries requires more computational resources to explore the unnecessary regions in the Hilbert space. The conventional Subspace-Search VQE (SSVQE) algorithm, which can calculate excited states, is also unable to effectively handle degenerate states since the loss function only contains the expectation value of the Hamiltonian. In this study, the energy eigenstates of the one-dimensional Fermi-Hubbard model with two lattice sites and the two-dimensional Hubbard model with four lattice sites are calculated. By incorporating symmetries into the quantum circuits and loss function, we find that both the ground state and excited state calculations are improved greatly compared to the case without symmetries. The enhancement in excited state calculations is particularly significant. This is because quantum circuits that conserve the particle number are used, and appropriate penalty terms are added to the loss function, enabling the optimization process to correctly identify degenerate states. The results are verified through repeated simulations.
@@ I INTRODUCTION
The concept of quantum computing was first proposed by Feynman [1] due to the difficulties in the large scale quantum system simulation conducted on the classical computers. Subsequently, Shor’s algorithm, proposed by Shor based on quantum computers[2], attracted widespread attention to quantum computing for the first time, as it could break the Rivest-Shamir-Adleman (RSA) cryptosystem in polynomial time. This was also the first significant application of quantum computing beyond quantum simulation. Since quantum algorithms were first truly executed on a two-qubit experimental device [3] and Google company claimed to achieve quantum supremacy [4], the impact of quantum computing has been growing up significantly. Currently, limited by technology, experimental devices equipped with qubits are referred to as noisy intermediate scale quantum (NISQ) information processors [5, 6, 7]. This is due to the fact that quantum computers cannot scale to a sufficient size. In other words, quantum computers cannot host a large enough number of qubits. Additionally, the presence of noise implies that the error effects must be taken into account.
@@
On current NISQ devices, in order to harness the advantages of quantum computing, the Variational Quantum Eigensolver (VQE) is proposed as a hybrid quantum-classical algorithm that is capable of effectively utilizing quantum computing [8, 9, 10, 11, 12, 13, 14, 15]. It combines the advantages of quantum and classical computing, respectively. VQE enables the implementation of quantum algorithms that have a clear advantage over classical algorithms when running on a quantum computer. The optimization processes, currently beyond the capabilities of quantum computers, are handled by classical computers. This collaboration between quantum and classical systems significantly enhances computational power. The loss function in VQE algorithm for the optimization is defined as the expectation value of the system’s Hamiltonian. The ground state of the system can be obtained by minimizing the loss function. Based on this algorithm, Nakanishi et al. further developed the Subspace-Search VQE (SSVQE) algorithm [16], of which the loss function is defined as a weighted sum of the Hamiltonian expectation values over $n$ orthogonal initial states. Then the first $n$ energy eigenstates can be calculated in case of non-degenerate states.
@@
For the quantum many-body systems, the dimension of the Hilbert space that describes the state increases exponentially along with the increase of the number of particles in the system. The dimension of the Hamiltonian matrix to be solved by classical computers also increases exponentially accordingly, making the computation blows up. This is also known as the “exponential wall” problem [17]. However, no matter VQE or SSVQE algorithm, the symmetries of the system, under which the unnecessary regions in Hilbert space can be skipped during the optimization process, are not taken into account. Nevertheless, they fail to effectively distinguish degenerate state in the calculation of excited states. This is because of that the SSVQE algorithm only includes terms related to the Hamiltonian expectation value. To resolve these issues, a symmetry-preserving method was developed by Lyu et. al [18] and verified on the experimental devices [19]. The hardware and hybrid symmetry-preserving methods were employed to calculate the energy eigenstates of the one-dimensional Heisenberg model. Note that symmetry is considered only in the quantum circuit in the first one, while symmetry is included in both the quantum circuit and the loss function in the second one.
@@
The Hubbard model is a model used in condensed matter physics to describe strongly correlated electron systems. Although the exact solution to the one-dimensional Hubbard model has long been found [20], the solutions for the two-dimensional or three-dimensional Hubbard models have not yet been found. As an important model for understanding the high-temperature superconducting mechanisms of materials like uprate  [21], solving the Hubbard model is of great significance. As for the quantum many-body system, the Fermi-Hubbard model [12, 13, 22, 23, 24, 25, 26] also faces the exponential wall problem along with the increase of the number of lattice sites. The Hubbard Hamiltonian can transfer from fermionic creation and annihilation operators into a finite-dimensional matrix form by Jordan-Wigner transformation [27, 28]. The energy eigenstates of the system can be obtained by diagonalizing the matrix. However, the matrix dimension increases as $4^{n}$ with the number of lattice sites $n$. e.g. a matrix of size $4^{15}\approx 10^{9}$ needs to be solved in case of a system with 15 lattice sites. This numerical calculation task far exceeds the capability of the most powerful supercomputers currently in the world which can only handle eigenvalue problems up to $\approx 10^{8}$ dimensions [29]. The details of the evaluation are shown in Appendix  A. In contrast, it is possible that just 30 qubits are needed to solve the problem with the quantum computers.
@@
To reduce the scale of the calculation, symmetry can be taken into account in VQE algorithm for the Hubbard model. Thus in this paper, the ground state of a one-dimensional Fermi-Hubbard model with two lattice sites and a two-dimensional Hubbard model with four lattice sites are chosen for the exploration of the VQE with symmetry. The SSVQE algorithm, with symmetries both in quantum circuit and the loss function, is applied to solve for the excited states. The effectiveness of the symmetry enhanced approach in solving the eigenstates of the Hubbard model will be checked in detail, which may give us a hint for more large many-body systems. In all, the paper is organized as following: the VQE algorithm is introduced in Sec. II. The Jordan-Wigner transformation is employed in Sec.  III to convert fermionic creation and annihilation operators into strings of Pauli matrices and to encode the Hubbard operators. The energy eigenstates of the system using the symmetry-enhanced VQE algorithm are studied in Sec. IV and sec. V. Finally, our conclusions is given in sec. VI.
@@ II.1 VQE for ground state
An optimization algorithm, such as gradient descent [30], is used to update the parameters. The updated parameters are then passed to the quantum circuit, and such process is repeated. After multiple iterations, the minimized loss function is obtained, $\mathcal{L}(\bm{\theta}^{*})=\underset{\bm{\theta}}{\text{min }}\mathcal{L}\left(\bm{\theta}\right)$, corresponding to the ground-state energy. The ground-state vector is obtained by applying the optimized quantum circuit to the initial state $|E_{0}\rangle=U(\bm{\theta}^{*})|\psi_{I}\rangle$.
For a single qubit, the variational parameters in the quantum circuit represent the rotation angles of the qubit’s state vector around the x, y, or z axes on the Bloch sphere [31]. As for a two-state system $|\psi\rangle=\alpha|0\rangle+\beta|1\rangle$, it can be transformed in the two-dimensional complex vector space by an element $U\left(a,b\right)$ from the SU(2) group, where $\left|a\right|^{2}+\left|b\right|^{2}=1$, resulting in $|\psi^{{}^{\prime}}\rangle=U|\psi\rangle$. This corresponds to a rotational transformation in three-dimensional real vector space. Since SU(2) group and the SO(3) group are homomorphic, We can denote that SO(3) group are represented by $a$, $b$ as $R\left(a,b\right)$ (the specific matrix representation is given in Appendix  B). Set $a=e^{-i\frac{\theta}{2}}$, $b=0$, then we have
$R\left(a=e^{-i\frac{\theta}{2}},b=0\right)=\begin{pmatrix}\cos\theta&-\sin\theta&0\\ \sin\theta&\cos\theta&0\\ 0&0&1\end{pmatrix}\,.$ (2)
This SO(3) element corresponds to a rotation around the z-axis. If $a=e^{-i\frac{\theta}{2}},b=0$ is substituted into the group element of the SU(2) group, we can get
$U\left(a^{{}^{\prime}},b^{{}^{\prime}}\right)$ $=$ $\begin{pmatrix}a&b\\ -b^{*}&a^{*}\end{pmatrix}\Bigg{|}_{a=e^{-i\frac{\theta}{2}},b=0}$ (3)
$=$ $\begin{pmatrix}e^{-i\frac{\theta}{2}}&0\\ 0&e^{i\frac{\theta}{2}}\end{pmatrix}\,,$
which corresponds to $R_{z}\left(\theta\right)=e^{-i\frac{\theta}{2}\sigma_{z}}$, a commonly used quantum gate. The variation of the group parameter $\theta$ corresponds to the rotation of the vector around the z-axis in three-dimensional space. Similarly, $R_{x}\left(\theta\right)=e^{-i\frac{\theta}{2}\sigma_{x}}$ and $R_{y}\left(\theta\right)=e^{-i\frac{\theta}{2}\sigma_{y}}$. This also implies that Pauli matrices $\sigma_{x}$, $\sigma_{y}$ and $\sigma_{z}$ are the generators of the state vector rotations around the x-axis, y-axis and z-axis, respectively.
@@ II.2 VQE for excited state
Although VQE is an effective algorithm, it is limited in calculating the ground state and cannot solve for excited states. Nakanishi et al. [16] developed a method, known as the Weighted SSVQE algorithm or simply the SSVQE algorithm, which can efficiently compute the first $k$ eigenstates of a system. This algorithm selects a series of mutually orthogonal initial states and optimizes a loss function defined as the weighted sum of the Hamiltonian’s expectation values for different initial states and weights, thereby obtaining the respective energy eigenstates.
@@
where $\omega_{0}>\omega_{1}>\cdots>\omega_{k-1}$. By optimizing the loss function to approach the minimum value of $\mathcal{L}(\bm{\theta}^{*})=\underset{\bm{\theta}}{\text{min }}\mathcal{L}\left(\bm{\theta}\right)$, we can obtain the energy eigenvalues
@@ III.1 The Jordan-Wigner transformation
The Jordan-Wigner transformation was initially developed to map the spin operators to fermionic creation and annihilation operators [27]. Later, for the purposes of computation and measurement, it became necessary to transform Hamiltonians expressed in terms of creation and annihilation operators into combinations of Pauli operators [28] in quantum computing. The process is exactly the inverse of the original transformation. The raising and lowering operators for qubit states are defined as $\sigma^{-}=|0\rangle\langle 1|=\frac{1}{2}\left(\sigma_{x}+i\sigma_{y}\right)$ and $\sigma^{+}=|1\rangle\langle 0|=\frac{1}{2}\left(\sigma_{x}-i\sigma_{y}\right)$, where $\sigma_{x}$ and $\sigma_{y}$ are Pauli matrices. In the following, the identity matrix and Pauli matrices are denoted as $I$, $X$, $Y$ and $Z$.
@@ III.2 Operators Encoding
where $\sigma\in\{\uparrow,\downarrow\}$ and $\hat{n}_{i\sigma}=\hat{a}_{i\sigma}^{\dagger}\hat{a}_{i\sigma}$. The term $\hat{a}_{i\sigma}^{\dagger}\hat{a}_{j\sigma}$ represents the annihilation of an electron with spin $\sigma$ at site $j$ and the creation of an electron with spin $\sigma$ at site $i$, describing electron hopping between sites $i$ and $j$. T he summation $\sum_{i,j}\sum_{\sigma}$ accounts for all the nearest-neighbor sites and both spin states. Here, $t$ is the hopping amplitude, indicating the probability of electron hopping between two lattice sites. The operator $\hat{n}_{i\sigma}=\hat{a}_{i\sigma}^{\dagger}\hat{a}_{i\sigma}$ is the particle number operator for electrons with spin $\sigma$. $\hat{n}_{i\uparrow}\hat{n}_{i\downarrow}$ denotes the number of electron pairs with opposite spins on site $i$ (taking values of 0 or 1). Finally, $U$ represents the Coulomb interaction potential between two electrons of opposite spin on the same lattice site. In this study, we choose $t=1$ and $U=2$. This is a common choice [32, 33].
@@
The Hubbard Hamiltonian commutes with the particle number operator $\hat{N}$ and the $z$-component of the total spin operator $\hat{S}_{z}$ in this system [34, 35]:
@@ IV.2.1 Ground state
Table 1: Eigenstates with corresponding energy, particle number, and $z$-component of the total spin for a two-site 1D Hubbard model.
@@
$|\alpha_{2,0}\rangle$ $-1.2361$ $2$ 00
$|\beta_{1,-\frac{1}{2}}\rangle$ $-1.0000$ $1$ $-\frac{1}{2}$
$|\beta_{1,\frac{1}{2}}\rangle$ $-1.0000$ $1$ $\frac{1}{2}$
$|\gamma_{2,-1}\rangle$ $0.0000$ $2$ $-1$
$|\gamma_{0,0}\rangle$ $0000.00$
$|\gamma_{2,1}\rangle$ $0.0000$ $2$ $1$
$|\gamma_{2,0}\rangle$ $0.0000$ $2$ 00
$|\delta_{3,-\frac{1}{2}}\rangle$ $1.0000$ $3$ $-\frac{1}{2}$
$|\delta_{1,\frac{1}{2}}\rangle$ $1.0000$ $1$ $\frac{1}{2}$
$|\delta_{1,-\frac{1}{2}}\rangle$ $1.0000$ $1$ $-\frac{1}{2}$
$|\delta_{3,\frac{1}{2}}\rangle$ $1.0000$ $3$ $\frac{1}{2}$
$|\varepsilon_{2,0}\rangle$ $2.0000$ $2$ 00
$|\zeta_{3,-\frac{1}{2}}\rangle$ $3.0000$ $3$ $-\frac{1}{2}$
$|\zeta_{3,\frac{1}{2}}\rangle$ $3.0000$ $3$ $\frac{1}{2}$
$|\eta_{2,0}\rangle$ $3.2361$ $2$ 00
$|\theta_{4,0}\rangle$ $4.0000$ $4$ 00
@@
When constructing the initial state, it is necessary to know the particle number and the $z$-component of spin corresponding to that state. Therefore, we need to determine the eigenvalue equations of these two operators in the computational basis. The following analysis provides the details. The two eigenstates of the qubit correspond to the encoding of the spin-up and spin-down occupation states and vacant states of electrons at each lattice site. Therefore, it is evident that each computational basis corresponds to the eigenstates of the $S_{z}$ operator. Qubits 0-1 encode the occupation states and vacant states of spin-up electrons. The contribution to the $S_{z}$ eigenvalue is 0 when the qubit state is $|0\rangle$ and $\frac{1}{2}$ when the qubit state is $|1\rangle$. Qubits 2-3 encode the occupation states and vacant states of spin-down electrons. The contribution to the $S_{z}$ eigenvalue is 0 when the qubit state is $|0\rangle$ and $-\frac{1}{2}$ when the qubit state is $|1\rangle$. For example, the eigenvalue of $\hat{S}_{z}$ operator for the state $|0_{0}1_{1}1_{2}1_{3}\rangle$ is $\left(\frac{1}{2}\right)_{1}+\left(-\frac{1}{2}\right)_{2}+\left(-\frac{1}{2}\right)_{3}=-\frac{1}{2}$, where the subscript indicates the qubit index.
@@
The first symmetry-preserving ansatz we use consists of two-qubit gates as the basic unit [36, 37],
$U_{P}\left(\bm{\theta}\right)=\begin{pmatrix}1&0&0&0\\ 0&\cos\theta_{1}&e^{i\theta_{2}}\sin\theta_{1}&0\\ 0&e^{-i\theta_{2}}\sin\theta_{1}&-\cos\theta_{1}&0\\ 0&0&0&1\end{pmatrix}\,.$ (20)
This two-qubit gate can preserve the particle number of states like $|01\rangle$ or $|10\rangle$. For example, $U_{P}\left(\bm{\theta}\right)|01\rangle=\cos\theta_{1}|01\rangle+e^{-i\theta_{2}}\sin\theta_{1}|10\rangle$. This is satisfied for the initial state we selected. However, this gate can only preserve the particle number for such states and does not preserve it for states like $|00\rangle$ or $|11\rangle$. This limitation means that this quantum gate cannot be used as a symmetry-preserving method to solve for other eigenstates. The quantum circuit corresponding to this gate is shown in Fig. 1(c). The quantum circuit for applying it to four qubits in a single layer is shown in Fig. 1(d). Two layers of this circuit are used to generate the ground state, containing 48 parameters and 36 CNOT gates. The first iteration where the average fidelity $\geq 0.99$ occurs at iteration 33, showing a noticeable improvement over the hardware efficient ansatz. This is because the symmetry-preserving ansatz keeps the particle number fixed, so we do not have to search the state space for states that do not match the required particle number. Figure 2(g) shows the expectation value of the particle number operator $\langle\psi_{0}\left(\bm{\theta}\right)|\hat{N}|\psi_{0}\left(\bm{\theta}\right)\rangle$, which remains constant throughout the computation. In contrast, for the hardware efficient ansatz, Fig. 2(c) requires optimization over a large state space to reach the ground state with the correct particle number.
The other symmetry-preserving ansatz we use consists of two-qubit gates as the basic unit [38],
$U_{N}\left(\theta\right)=e^{i\theta\left(\sigma_{x}\otimes\sigma_{x}+\sigma_{y}\otimes\sigma_{y}+\sigma_{z}\otimes\sigma_{z}\right)}\,.$ (21)

### v2

@@ Abstract
The Variational Quantum Eigensolver (VQE), as a hybrid quantum-classical algorithm, is an important tool for effective quantum computing in the current noisy intermediate-scale quantum (NISQ) era. However, the traditional hardware-efficient ansatz without taking into account symmetries requires more computational resources to explore the unnecessary regions in the Hilbert space. The conventional Subspace-Search VQE (SSVQE) algorithm, which can calculate excited states, is also unable to effectively handle degenerate states since the loss function only contains the expectation value of the Hamiltonian. In this study, the energy eigenstates of the one-dimensional Fermi-Hubbard model with two and four lattice sites are calculated, respectively. By incorporating symmetries into the quantum circuits and loss function, we find that both the ground state and excited state calculations are improved greatly compared to the case without symmetries. The enhancement in excited state calculations is particularly significant. This is because quantum circuits that conserve the particle number are used, and appropriate penalty terms are added to the loss function, enabling the optimization process to correctly identify degenerate states. The results are verified through repeated simulations.
Quantum simulation, Hubbard model, Variational quantum eigensolver, Quantum computation
@@ I INTRODUCTION
The concept of quantum computing was first proposed by Feynman [Feynman1982] due to the difficulties in the large scale quantum system simulation conducted on the classical computers. Subsequently, Shor’s algorithm, proposed by Shor based on quantum computers[Shor1994], attracted widespread attention to quantum computing for the first time, as it could break the Rivest-Shamir-Adleman (RSA) cryptosystem in polynomial time. This was also the first significant application of quantum computing beyond quantum simulation. Since quantum algorithms were first truly executed on a two-qubit experimental device [Chuang1998] and Google company claimed to achieve quantum supremacy [Google2019], the impact of quantum computing has been growing up significantly. Currently, limited by technology, experimental devices equipped with qubits are referred to as noisy intermediate scale quantum (NISQ) information processors [NISQ2018, NISQMaterials2021, NISQChemistry2019]. This is due to the fact that quantum computers cannot scale to a sufficient size. In other words, quantum computers cannot host a large enough number of qubits. Additionally, the presence of noise implies that the error effects must be taken into account.
@@
On current NISQ devices, in order to harness the advantages of quantum computing, the Variational Quantum Eigensolver (VQE) is proposed as a hybrid quantum-classical algorithm that is capable of effectively utilizing quantum computing [VQEOriginal2014, VQEOriginal2016, VQEReview2022, VQESimulation2023, VQEHubbard2022, VQEDMETHubbard2022, VQEMolecular2016, VQEMapping2023]. It combines the advantages of quantum and classical computing, respectively. VQE enables the implementation of quantum algorithms that have a clear advantage over classical algorithms when running on a quantum computer. The optimization processes, currently beyond the capabilities of quantum computers, are handled by classical computers. This collaboration between quantum and classical systems significantly enhances computational power. The loss function in VQE algorithm for the optimization is defined as the expectation value of the system’s Hamiltonian. The ground state of the system can be obtained by minimizing the loss function. Based on this algorithm, Nakanishi et al. further developed the Subspace-Search VQE (SSVQE) algorithm [SSVQE2019], of which the loss function is defined as a weighted sum of the Hamiltonian expectation values over $n$ orthogonal initial states. Then the first $n$ energy eigenstates can be calculated in case of non-degenerate states.
@@
For the quantum many-body systems, the dimension of the Hilbert space that describes the state increases exponentially along with the increase of the number of particles in the system. The dimension of the Hamiltonian matrix to be solved by classical computers also increases exponentially accordingly, making the computation blows up. This is also known as the “exponential wall” problem [ExponentialWall2010]. However, no matter VQE or SSVQE algorithm, the symmetries of the system, under which the unnecessary regions in Hilbert space can be skipped during the optimization process, are not taken into account. Nevertheless, they fail to effectively distinguish degenerate state in the calculation of excited states. This is because of that the SSVQE algorithm only includes terms related to the Hamiltonian expectation value. To resolve these issues, a symmetry-preserving method was developed by Lyu et. al [SymmetryEnhancedVQE2023] and verified on the experimental devices [SymmetryEnhancedVQE2024]. The hardware and hybrid symmetry-preserving methods were employed to calculate the energy eigenstates of the one-dimensional Heisenberg model. Note that symmetry is considered only in the quantum circuit in the first one, while symmetry is included in both the quantum circuit and the loss function in the second one.
@@
The Hubbard model is a model used in condensed matter physics to describe strongly correlated electron systems. Although the exact solution to the one-dimensional Hubbard model has long been found [1DHubbard1992], the solutions for the two-dimensional or three-dimensional Hubbard models have not yet been found. As an important model for understanding the high-temperature superconducting mechanisms of materials like uprate  [CuprateHubbard2024], solving the Hubbard model is of great significance. As for the quantum many-body system, the Fermi-Hubbard model [VQEHubbard2022, VQEDMETHubbard2022, HubbardModel2022, PanJianweiHubbard2024, LowDepthHubbard2019, TetrahedronHubbard1995, MonteCarloHubbard1997] also faces the exponential wall problem along with the increase of the number of lattice sites. The Hubbard Hamiltonian can transfer from fermionic creation and annihilation operators into a finite-dimensional matrix form by Jordan-Wigner transformation [JordanWigner1928, JordanWigner2011]. The energy eigenstates of the system can be obtained by diagonalizing the matrix. However, the matrix dimension increases as $4^{n}$ with the number of lattice sites $n$. e.g. a matrix of size $4^{15}\approx 10^{9}$ needs to be solved in case of a system with 15 lattice sites. This numerical calculation task far exceeds the capability of the most powerful supercomputers currently in the world which can only handle eigenvalue problems up to $\approx 10^{8}$ dimensions [Supercomputer2023]. The details of the evaluation are shown in Appendix  A. In contrast, it is possible that just 30 qubits are needed to solve the problem with the quantum computers.
@@
To reduce the scale of the calculation, symmetry can be taken into account in VQE algorithm for the Hubbard model. In this study, regarding the quantum circuit, the invariance of the number of qubit excitations proposed in [ParticleCircuit2020, ParticleCircuit2023, ParticleCircuit2004] is connected to the particle number symmetry of the Hubbard system and applied to realistic and physically meaningful systems. Moreover, symmetry is considered not only within the quantum circuit but also in the loss function, and is further utilized to solve degenerate excited states. Previous studies on solving the Hubbard model mostly focused on ground state solutions and did not emphasize the system’s symmetry. In all, the paper is organized as following: the VQE algorithm is introduced in Sec. II. The Jordan-Wigner transformation is employed in Sec.  III to convert fermionic creation and annihilation operators into strings of Pauli matrices and to encode the Hubbard operators. The energy eigenstates of the system using the symmetry-enhanced VQE algorithm are studied in Sec. IV and sec. V. Finally, our conclusions is given in sec. VI.
@@ II.1 VQE for ground state
[…]
For a single qubit, the variational parameters in the quantum circuit represent the rotation angles of the qubit’s state vector around the $x$, $y$, or $z$ axes on the Bloch sphere [BlochSphere2019]. As for a two-state system $|\psi\rangle=\alpha|0\rangle+\beta|1\rangle$, it can be transformed in the two-dimensional complex vector space by an element $U\left(a,b\right)$ from the SU(2) group, where $\left|a\right|^{2}+\left|b\right|^{2}=1$, resulting in $|\psi^{{}^{\prime}}\rangle=U|\psi\rangle$. This corresponds to a rotational transformation in three-dimensional real vector space. Since SU(2) group and the SO(3) group are homomorphic, We can denote that SO(3) group are represented by $a$, $b$ as $R\left(a,b\right)$ (the specific matrix representation is given in Appendix  B). Set $a=\mathrm{e}^{-\mathrm{i}\frac{\theta}{2}}$, $b=0$, then we have
$R\left(a=\mathrm{e}^{-\mathrm{i}\frac{\theta}{2}},b=0\right)=\begin{pmatrix}\cos\theta&-\sin\theta&0\\ \sin\theta&\cos\theta&0\\ 0&0&1\end{pmatrix}\,$. (2)
This SO(3) element corresponds to a rotation around the $z$-axis. If $a=\mathrm{e}^{-\mathrm{i}\frac{\theta}{2}},b=0$ is substituted into the group element of the SU(2) group, we can get
$U\left(a^{{}^{\prime}},b^{{}^{\prime}}\right)$ $=$ $\begin{pmatrix}a&b\\ -b^{*}&a^{*}\end{pmatrix}\Bigg|_{a=\mathrm{e}^{-\mathrm{i}\frac{\theta}{2}},b=0}$ (3)
$=$ $\begin{pmatrix}\mathrm{e}^{-\mathrm{i}\frac{\theta}{2}}&0\\ 0&\mathrm{e}^{\mathrm{i}\frac{\theta}{2}}\end{pmatrix}\,$,
which corresponds to $R_{z}\left(\theta\right)=\mathrm{e}^{-\mathrm{i}\frac{\theta}{2}\sigma_{z}}$, a commonly used quantum gate. The variation of the group parameter $\theta$ corresponds to the rotation of the vector around the $z$-axis in three-dimensional space. Similarly, $R_{x}\left(\theta\right)=\mathrm{e}^{-\mathrm{i}\frac{\theta}{2}\sigma_{x}}$ and $R_{y}\left(\theta\right)=\mathrm{e}^{-\mathrm{i}\frac{\theta}{2}\sigma_{y}}$. This also implies that Pauli matrices $\sigma_{x}$, $\sigma_{y}$ and $\sigma_{z}$ are the generators of the state vector rotations around the $x$-axis, $y$-axis and $z$-axis, respectively.
@@ II.2 VQE for excited state
Although VQE is an effective algorithm, it is limited in calculating the ground state and cannot solve for excited states. Nakanishi et al. [SSVQE2019] developed a method, known as the Weighted SSVQE algorithm or simply the SSVQE algorithm, which can efficiently compute the first $k$ eigenstates of a system. This algorithm selects a series of mutually orthogonal initial states and optimizes a loss function defined as the weighted sum of the Hamiltonian’s expectation values for different initial states and weights, thereby obtaining the respective energy eigenstates.
@@
[…]
@@ III.1 The Jordan-Wigner transformation
The Jordan-Wigner transformation was initially developed to map the spin operators to fermionic creation and annihilation operators [JordanWigner1928]. Later, for the purposes of computation and measurement, it became necessary to transform Hamiltonians expressed in terms of creation and annihilation operators into combinations of Pauli operators [JordanWigner2011] in quantum computing. The process is exactly the inverse of the original transformation. The raising and lowering operators for qubit states are defined as $\sigma^{-}=|0\rangle\langle 1|=\frac{1}{2}\left(\sigma_{x}+\mathrm{i}\sigma_{y}\right)$ and $\sigma^{+}=|1\rangle\langle 0|=\frac{1}{2}\left(\sigma_{x}-\mathrm{i}\sigma_{y}\right)$, where $\sigma_{x}$ and $\sigma_{y}$ are Pauli matrices. In the following, the identity matrix and Pauli matrices are denoted as $I$, $X$, $Y$ and $Z$.
@@ III.2 Operators Encoding
where $\sigma\in\{\uparrow,\downarrow\}$ and $\hat{n}_{i\sigma}=\hat{a}_{i\sigma}^{\dagger}\hat{a}_{i\sigma}$. The term $\hat{a}_{i\sigma}^{\dagger}\hat{a}_{j\sigma}$ represents the annihilation of an electron with spin $\sigma$ at site $j$ and the creation of an electron with spin $\sigma$ at site $i$, describing electron hopping between sites $i$ and $j$. T he summation $\sum_{i,j}\sum_{\sigma}$ accounts for all the nearest-neighbor sites and both spin states. Here, $t$ is the hopping amplitude, indicating the probability of electron hopping between two lattice sites. The operator $\hat{n}_{i\sigma}=\hat{a}_{i\sigma}^{\dagger}\hat{a}_{i\sigma}$ is the particle number operator for electrons with spin $\sigma$. $\hat{n}_{i\uparrow}\hat{n}_{i\downarrow}$ denotes the number of electron pairs with opposite spins on site $i$ (taking values of 0 or 1). Finally, $U$ represents the Coulomb interaction potential between two electrons of opposite spin on the same lattice site. In this study, we choose $t=1$ and $U=2$. This is a common choice [HubbardtU2015, HubbardtU2020].
@@
The Hubbard Hamiltonian commutes with the particle number operator $\hat{N}$ and the $z$-component of the total spin operator $\hat{S}_{z}$ in this system [SymmetryHubbard2022, SymmetryHubbard1989]:
@@ IV.2.1 Ground state
Table 1: Eigenstates with corresponding energy, particle number, and $z$-component of the total spin for a two-site Hubbard model.
@@
$|\alpha_{2,0}\rangle$ 1.2361-1.2361 22 0
$|\beta_{1,-\frac{1}{2}}\rangle$ 1.0000-1.0000 11 $\frac{1}{2}$
$|\beta_{1,\frac{1}{2}}\rangle$ 1.0000-1.0000 11 $\frac{1}{2}$
$|\gamma_{2,-1}\rangle$ $0.0000$ 22 1-1
$|\gamma_{0,0}\rangle$ $0.0000$ 0 0
$|\gamma_{2,1}\rangle$ $0.0000$ 22 11
$|\gamma_{2,0}\rangle$ $0.0000$ 22 0
$|\delta_{3,-\frac{1}{2}}\rangle$ $1.0000$ 33 $\frac{1}{2}$
$|\delta_{1,\frac{1}{2}}\rangle$ $1.0000$ 11 $\frac{1}{2}$
$|\delta_{1,-\frac{1}{2}}\rangle$ $1.0000$ 11 $\frac{1}{2}$
$|\delta_{3,\frac{1}{2}}\rangle$ $1.0000$ 33 $\frac{1}{2}$
$|\varepsilon_{2,0}\rangle$ $2.0000$ 22 0
$|\zeta_{3,-\frac{1}{2}}\rangle$ $3.0000$ 33 $\frac{1}{2}$
$|\zeta_{3,\frac{1}{2}}\rangle$ $3.0000$ 33 $\frac{1}{2}$
$|\eta_{2,0}\rangle$ $3.2361$ 22 0
$|\theta_{4,0}\rangle$ $4.0000$ 44 0
@@
When constructing the initial state, it is necessary to know the particle number and the $z$-component of spin corresponding to that state. Therefore, we need to determine the eigenvalue equations of these two operators in the computational basis. The following analysis provides the details. The two eigenstates of the qubit correspond to the encoding of the spin-up and spin-down occupation states and vacant states of electrons at each lattice site. Therefore, it is evident that each computational basis corresponds to the eigenstates of the $S_{z}$ operator. Qubits 0-1 encode the occupation states and vacant states of spin-up electrons. The contribution to the $S_{z}$ eigenvalue is 0 when the qubit state is $|0\rangle$ and $\frac{1}{2}$ when the qubit state is $|1\rangle$. Qubits 2-3 encode the occupation states and vacant states of spin-down electrons. The contribution to the $S_{z}$ eigenvalue is 0 when the qubit state is $|0\rangle$ and $\frac{1}{2}$ when the qubit state is $|1\rangle$. For example, the eigenvalue of $\hat{S}_{z}$ operator for the state $|0_{0}1_{1}1_{2}1_{3}\rangle$ is $\left(\frac{1}{2}\right)_{1}+\left(-\frac{1}{2}\right)_{2}+\left(-\frac{1}{2}\right)_{3}=-\frac{1}{2}$, where the subscript indicates the qubit index.
@@
The first symmetry-preserving ansatz we use consists of two-qubit gates as the basic unit [ParticleCircuit2020, ParticleCircuit2023],
$U_{P}\left(\bm{\theta}\right)=\begin{pmatrix}1&0&0&0\\ 0&\cos\theta_{1}&\mathrm{e}^{\mathrm{i}\theta_{2}}\sin\theta_{1}&0\\ 0&\mathrm{e}^{-\mathrm{i}\theta_{2}}\sin\theta_{1}&-\cos\theta_{1}&0\\ 0&0&0&1\end{pmatrix}\,$. (20)
This two-qubit gate can preserve the particle number of states like $|01\rangle$ or $|10\rangle$. For example, $U_{P}\left(\bm{\theta}\right)|01\rangle=\cos\theta_{1}|01\rangle+\mathrm{e}^{-\mathrm{i}\theta_{2}}\sin\theta_{1}|10\rangle$. This is satisfied for the initial state we selected. However, this gate can only preserve the particle number for such states and does not preserve it for states like $|00\rangle$ or $|11\rangle$. This limitation means that this quantum gate cannot be used as a symmetry-preserving method to solve for other eigenstates. The quantum circuit corresponding to this gate is shown in Fig. 1(c). The quantum circuit for applying it to four qubits in a single layer is shown in Fig. 1(d). Two layers of this circuit are used to generate the ground state, containing 48 parameters and 36 CNOT gates. The first iteration where the average fidelity $\geq 0.99$ occurs at iteration 33, showing a noticeable improvement over the hardware efficient ansatz. This is because the symmetry-preserving ansatz keeps the particle number fixed, so we do not have to search the state space for states that do not match the required particle number. Figure 2(g) shows the expectation value of the particle number operator $\langle\psi_{0}\left(\bm{\theta}\right)|\hat{N}|\psi_{0}\left(\bm{\theta}\right)\rangle$, which remains constant throughout the computation. In contrast, for the hardware efficient ansatz, Fig. 2(c) requires optimization over a large state space to reach the ground state with the correct particle number.
The other symmetry-preserving ansatz we use consists of two-qubit gates as the basic unit [ParticleCircuit2004],
$U_{N}\left(\theta\right)=\mathrm{e}^{\mathrm{i}\theta\left(\sigma_{x}\otimes\sigma_{x}+\sigma_{y}\otimes\sigma_{y}+\sigma_{z}\otimes\sigma_{z}\right)}\,.$ (21)